OpenStudy (anonymous):

Please help.... -1/9 + 4/9 =? -5/12 + 7/12 -4/5 - 3/5 = ? 1/12 - 7/12 = ? 1/4 + (-7/12) = ? -6/7 + (-1/2) = ? 1/3 - 7/8 = ? -2/9 - (-3/11) = ?

OpenStudy (anonymous):

you want ans?

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

"Give Help, Not Answers"

OpenStudy (anonymous):

i shall give you the ans if i had any but you can do these using the calculator much easier

OpenStudy (anonymous):

these are fraction right?

OpenStudy (anonymous):

i might be reading them wrong?

OpenStudy (littlebird):

What are you having difficulty with?

OpenStudy (anonymous):

does negative mins a negative turn into a postive with fractions too?

OpenStudy (anonymous):

minus*

OpenStudy (anonymous):

no

OpenStudy (anonymous):

only when its multiplied or divided

OpenStudy (anonymous):

-2/-2 =2/2

OpenStudy (anonymous):

-2x-2 =4

OpenStudy (anonymous):

good luck!

OpenStudy (anonymous):

for the last few questions the b

OpenStudy (anonymous):

how do you get the same denominator

OpenStudy (littlebird):

-5 - -1= -4 I just subtracted a negative from a negative and got a negative. -5 - -9= 4 I just subtracted a negative from a negative and got a positive.

OpenStudy (littlebird):

Here is an example of getting the same denominator

OpenStudy (anonymous):

thankyou :)

OpenStudy (anonymous):

-2/9 - (-3/11) = ? Yes, the double negative in the middle turns into a positive.